# atom sep=2em

### v1

@@ Abstract
Analyzing synthesis pathways for target molecules in a chemical reaction network annotated with information on the kinetics of individual reactions is an area of active study. This work presents a computational methodology for searching for pathways in reaction networks which is based on integer linear programming and the modeling of reaction networks by directed hypergraphs. Often multiple pathways fit the given search criteria. To rank them, we develop an objective function based on physical arguments maximizing the probability of the pathway. We furthermore develop an automated pipeline to estimate the energy barriers of individual reactions in reaction networks. Combined, the methodology facilitates flexible and kinetically informed pathway investigations on large reaction networks by computational means, even for networks coming without kinetic annotation, such as those created via generative approaches for expanding molecular spaces.
@@ I Introduction
Kinetic studies of individual reactions have provided valuable insight into their mechanisms, which in turn can generate ways to optimize the reaction to suit the overall synthesis goals. However, molecules are often entangled in various concurrent reactions, which together form an interconnected web of reactions called a reaction network. Theoretical and computational modeling offers a possibility for analyzing these systems to various degrees of accuracy and computational efficiency, depending on the level of abstraction employed. In this work, we propose a method for integrating kinetics into the search for pathways in reaction networks. The work is a continuation of our previous work [1], which focused on the thermodynamics of individual reactions in the network. However, despite the name, thermodynamics does not help to elucidate the ‘dynamics’ of the system: how fast does the pathway in question run and how quickly does the system equilibrate? For example, under atmospheric conditions, the conversion of diamond to graphite is thermodynamically favorable. However, the energy barrier for this conversion is high, so, thankfully, for all practical purposes ‘diamonds are forever’. Therefore, kinetic parameters of reactions, such as the free energy barrier or the rate constant, might provide better alternatives for characterizing pathways.
However, before one can use kinetic parameters to search for pathways in a reaction network, the network itself has to be constructed. This can be done in various ways. In one approach, the potential energy surface of a set of molecules is explored to elucidate trajectories with low-lying saddle points, separating the reactants from the products. Computationally intensive quantum mechanical calculations are used in most cases [2, 3], along with faster but less accurate semi-empirical methods in some [4, 5]. Recently, generative diffusion models have also been used to perturb molecules to generate reaction networks with unknown mechanisms [6, 7]. A different approach is to use reaction templates for iteratively expanding the network from a set of starting molecules, possibly in conjunction with filters to exclude specific reactions [8, 9, 10]. Further approaches include the generation of the reaction network of interest in a one-shot heuristical process by an expert [11, 12, 13], or the loading of an existing network from a database [14, 15].
@@
In this work, we develop an objective measure for ranking pathways in a reaction network designed to maximize the probability of the entire pathway based on physical principles. Additionally, based on the notion that energy barriers of reactions are the determining factor for the kinetics of the network, we develop an automated methodology for computing such energy barriers, building on a combination of an array of existing computational tools. A key novelty is that in contrast to the standard manual intervention required to generate reaction trajectories from which the reaction barrier can be estimated, we try to automate this part of the process. Lastly, we formulate an integer linear program which allows the user to specify queries for pathways in the network in a very flexible and generic way. Solving this integer linear program returns all pathways fulfilling the criteria, ranked by the objective measure mentioned above. In short, this gives an automated, general, and flexible methodology for analyzing reaction networks, a methodology which offers a speed-up over manual search for pathways but with comparable quality.
@@ III.1 From Rate Constants to Edge Weights
Since our modeling is not aimed at capturing temporal changes in the concentrations of molecules, we follow [30] in making the simplifying assumption that all molecules in the reaction network are present with unit concentrations. Under this assumption, the rate constant $k_{e}$ can be interpreted as being proportional to the probability of reaction $e$ happening in a fixed small volume and time interval. For a given reaction network normalize these values to sum to one:
@@ III.2 Formulating an Objective Function
$\max\left(-\sum_{e\in E}f_{e}\cdot\left(G_{e}-RT\cdot\log D\right)\right)$
@@
We note that via $D$, the objective value (4) of a pathway will depend on the underlying reaction network in which it is considered embedded. This may be regarded as a weakness. On the other hand, the objective value (4) is intuitively appealing by being composed of two natural components: Its first component $\sum_{e\in E}f_{e}\cdot G_{e}$ favors pathways with small sums of energy barriers of the reactions in the pathway, weighted by the number of times it is used. Its second component $\sum_{e\in E}f_{e}\cdot\left(RT\,\log D\right)$ favors pathways with fewer reactions. The factor $RT\log D$ can be viewed as weighting the relative importance of the two components.
@@ III.4.2 Method for Estimating Energy Barriers for the Reactions
  • For the energy barrier for the reaction to be estimated correctly, the order of the atoms in files containing the coordinates of the atoms for the endpoints must be the same. Our method of generating the reaction network (see Section LABEL:sec:_results) allows us to map corresponding atoms when applying a reaction template to a substructure match in a molecular graph [53]. This atom map is used to number the atoms on the two sides of the reaction correspondingly.
@@
IV Discussions
In this submission, we have described a methodology for kinetically informed exploration of pathways in reaction networks.
@@ IV Discussions
One part of the methodology is an automated workflow for annotating individual reactions with their energy barriers, computed using a diverse selection of tools— RDKit[63], xTB[47], ASE [55] and NeuralNEB [56]. The goal of this workflow is to minimize the laborious manual interventions and intensive expert knowledge often necessary when executing quantum chemical calculations to determine energy barriers. The workflow can be applied on networks of rather large scale, can work on the fly with generative network construction, and uses relatively moderate computational resources. The method does not take into account the interactions of solvent molecules while calculating the energy barriers. This and the question of how well models based on machine learning generalize past their training sets may lead to reasonable reservations on the precision of the estimates of the energy barriers for some use cases. However, our methodology offers the flexibility that this component can be substituted with alternatives, if desired.
@@
As the other part of the proposed methodology, we formulated an ILP to query pathways in the annotated network and designed an objective function for pathway evaluation. The query formalism is at the same time both precise and very flexible in what queries can be formulated. The objective function is made under the simplifying assumption (borrowed from [30]) that all molecules in the reaction network are present with unit concentrations at all times. Additionally, the costs assigned to pathways depend on the underlying reaction network considered, so comparisons across networks are not directly possible. However, the question of how to evaluate pathways is complex and the reader of course has the option to substitute a different objective function, based on other cost metrics. We do note, though, that our simplification and the formulation via ILP allows us to analyze much larger networks and using much less computational resources than alternatives like stochastic simulations or potential energy surface exploration. We also note that by a progressive constraining of the ILP, we in the end can return a list of high-scoring solutions, each representing structurally different pathways. Giving a list of good, structurally different solutions, rather than a single solution, helps mitigate disagreements about the details of pathway cost metrics. In summary, our work contributes to field of computationally assisted analysis and design of pathways in reaction networks.

### v2

@@ Abstract
Analyzing synthesis pathways for target molecules in a chemical reaction network annotated with information on the kinetics of individual reactions is an area of active study. This work presents a computational methodology for searching for pathways in reaction networks which is based on integer linear programming and the modeling of reaction networks by directed hypergraphs. Often multiple pathways fit the given search criteria. To rank them, we develop an objective function based on physical arguments maximizing the probability of the pathway. We furthermore develop an automated pipeline to estimate the energy barriers of individual reactions in reaction networks. Combined, the methodology facilitates flexible and kinetically informed pathway investigations on large reaction networks by computational means, even for networks coming without kinetic annotation, such as those created via generative approaches for expanding molecular spaces. To demonstrate the methodology, we apply it on a chemical reaction network modeling the molecular space of glycolonitrile chemistry, where we search for pathways to glycine and glycolic acid using small molecules as precursors.
@@ I Introduction
Kinetic studies of individual reactions have provided valuable insight into their mechanisms, which in turn can generate ways to optimize the reaction to suit the overall synthesis goals. However, molecules are often entangled in various concurrent reactions, which together form an interconnected web of reactions called a reaction network. Theoretical and computational modeling offers a possibility for analyzing these systems to various degrees of accuracy and computational efficiency, depending on the level of abstraction employed. In this work, we propose a method for integrating kinetics into the search for pathways in reaction networks modeled on a level of abstraction where the network is represented by a hypergraph and individual reactions are represented by directed hyperedges. The nodes of the hypergraph are molecular graphs (with nodes representing atoms and edges representing bonds) and hence the modeling is atomistically explicit. The work is a continuation of our previous work [1], which focused on the thermodynamics of individual reactions in the network. However, despite the name, thermodynamics does not help to elucidate the ‘dynamics’ of the system: how fast does the pathway in question run and how quickly does the system equilibrate? For example, under atmospheric conditions, the conversion of diamond to graphite is thermodynamically favorable. However, the energy barrier for this conversion is high, so, thankfully, for all practical purposes ‘diamonds are forever’. Therefore, kinetic parameters of reactions, such as the free energy barrier or the rate constant, might provide better alternatives for characterizing pathways.
However, before one can use kinetic parameters to search for pathways in a reaction network, the network itself has to be constructed. This can be done in various ways. In one approach, the potential energy surface of a set of molecules is explored to elucidate trajectories with low-lying saddle points, separating the reactants from the products. Computationally intensive quantum mechanical calculations are used in most cases [2, 3, 4], along with faster but less accurate semi-empirical methods in some [5, 6]. Recently, generative diffusion models have also been used to perturb molecules to generate reaction networks with unknown mechanisms [7, 8]. A different approach is to use reaction templates for iteratively expanding the network from a set of starting molecules, possibly in conjunction with filters to exclude specific reactions [9, 10, 11]. Further approaches include the generation of the reaction network of interest in a one-shot heuristical process by an expert [12, 13, 14], or the loading of an existing network from a database [15, 16].
@@
In this work, we develop an objective measure to rank pathways in a reaction network maximizing the probability of the entire pathway based on physical principles. Additionally, based on the notion that energy barriers of reactions are the determining factor for the kinetics of the network, we develop an automated methodology for computing such energy barriers, building on a combination of an array of existing computational tools. A key novelty is that in contrast to the standard manual intervention required to generate reaction trajectories from which the reaction barrier can be estimated, we try to automate this part of the process. Lastly, we formulate an integer linear program which allows the user to specify queries for pathways in the network in a very flexible and generic way. Solving this integer linear program returns all pathways fulfilling the criteria, ranked by the objective measure mentioned above. In short, this gives an automated, general, and flexible methodology for analyzing reaction networks, a methodology which offers a speed-up over manual search for pathways but with comparable quality. To exemplify its utility, we apply it to a reaction network expanded from the molecules glycolonitrile, ammonia and water, where we query for pathways to glycine, an amino acid, and glycolic acid, a molecule encountered in photorespiration in plants.
@@ III.1 From Rate Constants to Edge Weights
Since our modeling is not aimed at capturing temporal changes in the concentrations of molecules, we follow [31] in making the simplifying assumption that all molecules in the reaction network are present with unit concentrations. Under this assumption, the rate constant $k_{e}$ can be interpreted as being proportional to the probability of reaction $e$ happening in a fixed small volume and time interval. Also following [31], we for a given reaction network normalize these values to sum to one:
@@ III.2 Formulating an Objective Function
$\max\left(-\sum_{e\in E}f_{e}\cdot\left(G_{e}+RT\cdot\log D\right)\right)$
@@
We note that via $D$, the objective value (4) of a pathway will depend on the underlying reaction network in which it is considered embedded. This may be regarded as a weakness (also shared by [31]). On the other hand, the objective value (4) is intuitively appealing by being composed of two natural components: Its first component $\sum_{e\in E}f_{e}\cdot G_{e}$ favors pathways with small sums of energy barriers of the reactions in the pathway, weighted by the number of times it is used. Its second component $\sum_{e\in E}f_{e}\cdot\left(RT\,\log D\right)$ favors pathways with fewer reactions. The factor $RT\log D$ can be viewed as weighting the relative importance of the two components.
@@ III.4.2 Method for Estimating Energy Barriers for the Reactions
  • For the energy barrier for the reaction to be estimated correctly, the order of the atoms in files containing the coordinates of the atoms for the endpoints must be the same. Our method of generating the reaction network (see Section IV) allows us to map corresponding atoms when applying a reaction template to a substructure match in a molecular graph [54]. This atom map is used to number the atoms on the two sides of the reaction correspondingly.
@@
IV Results
In this section, we apply the methods from Section III to a reaction network involving glycolonitrile (MeCN), ammonia (NH3), and water (H2O), a network previously studied in [60]. The network was expanded using the graph transformation based generative framework MØD, starting from the three above-mentioned molecules, through recursive application of the graph transformation rules detailed in Appendix A.
Molecules with up to two carbon, four nitrogen, and four oxygen atoms were included in the generated network for comparability with [60]. The constructed reaction network had $44$ vertices and $116$ hyperedges, and shown in Figure 7. A list of molecules and reactions in the generated network is available in the GitHub data repository [data_repository]. To find probable pathways to glycine and glycolic acid, using small molecules MeCN (a hypothesized precursor to biotic molecules [61]), NH3, and H2O, an ILP was formulated to query the desired pathways. The ILP had $336$ variables and $1561$ constraints.
Figure 2: Top: A schematic depiction of the five best pathways to glyoxlic acid according to the chosen scoring function (energy levels not shown exactly to scale). The structures for the molecules in the corresponding vertices is listed in Table 3 and the weights assigned to the hyperedges listed in Table 4. Bottom: The energy profile of the best-scoring pathway (in blue) from above (energy levels shown to scale).
The five best scoring pathways to the target glycine is depicted in Figure 2. The best scoring pathway scores $143118$ on the chosen objective (III.2) and proceeds through iminoacetaldehyde ($v_{6}$), a pre-biotic intermediate hypothesized to be of significance [62, 63]. A water molecule acts as a catalyst in this pathway since it is used in reaction $e_{1}$ and generated back in reaction $e_{4}$. The second best pathway scores $143190$ on the objective (III.2). The last two reactions are the same for both (in fact, for all five) pathways. This second best pathway uses an ammonia molecule as a catalyst, molecules with a higher chemical potential (under unit concentrations) and reactions with higher energy barriers. The sum of the energy barriers for the reactions in this pathway is $200.02$ kJ mol-1 as opposed to $128.74$ kJ mol-1 for the best pathway. The third pathway branches off the second pathway and merges into the first, with a different reaction $e_{4}$. The fourth pathway uses a different higher energy molecule instead of iminoacetaldehyde. The fifth pathway has eight reactions, while the previous four had seven reactions. Despite having a lower sum of the energy barriers ($152.90$ kJ mol-1), it is ranked lower than the previous three pathways because it is longer. Figure 2 shows how interconnected the individual different pathways in the network might be, often with parallel branches and cross-connections, which leads to complex behavior of the system.
Figure 3: The six best pathways to glyoxlic acid according to the chosen scoring function. The structures for the molecules for the vertices is listed in Table 3 and the weights assigned to the hyperedges listed in Table 5.
Figure 3 shows the six best pathways to glycolic acid. The best scoring pathway has three reactions, the second best has four reaction, while the rest involve five reactions. Despite the third-listed pathway having a lower sum of energy barriers ($108.63$ kJ mol-1), it is ranked lower because it is longer, illustrating the effect of the second term in the objective (4).
V Discussions
@@
In this submission, we have described a methodology for kinetically informed exploration of pathways in reaction networks. One part of the methodology is an automated workflow for annotating individual reactions with their energy barriers, computed using a diverse selection of tools— RDKit[64], xTB[48], ASE [56] and NeuralNEB [57]. The goal of this workflow is to minimize the laborious manual interventions and intensive expert knowledge often necessary when executing quantum chemical calculations to determine energy barriers. As demonstrated in Section IV, the workflow can be applied on networks of rather large scale, can work on the fly with generative network construction, and uses relatively moderate computational resources. The method does not take into account the interactions of solvent molecules while calculating the energy barriers. This and the question of how well models based on machine learning generalize past their training sets may lead to reasonable reservations on the precision of the estimates of the energy barriers for some use cases. However, our methodology offers the flexibility that this component can be substituted with alternatives, if desired.
@@
As the other part of the proposed methodology, we formulated an ILP to query pathways in the annotated network and designed an objective function for pathway evaluation. The query formalism is at the same time both precise and very flexible in what queries can be formulated. The objective function is made under the simplifying assumption (borrowed from [31]) that all molecules in the reaction network are present with unit concentrations at all times. Additionally, the costs assigned to pathways depend on the underlying reaction network considered, so comparisons across networks are not directly possible. However, the question of how to evaluate pathways is complex [65, 66, 67] and the reader of course has the option to substitute a different objective function, based on other cost metrics. We do note, though, that our simplification and the formulation via ILP allows us to analyze much larger networks and using much less computational resources than alternatives like stochastic simulations or potential energy surface exploration. We also note that by a progressive constraining of the ILP, we in the end can return a list of high-scoring solutions, each representing structurally different pathways. Giving a list of good, structurally different solutions, rather than a single solution, helps mitigate disagreements about the details of pathway cost metrics.
Finally, we showcased the efficiency and utility of the described method on an actual reaction network, studying pathways to pre-biotically significant molecules, such as glycine and glycolic acid.
In summary, our work contributes to field of computationally assisted analysis and design of pathways in reaction networks represented as directed hypergraphs.
@@ Acknowledgements.
Appendix A Graph Transformation rules used
The graph transformation rules used to expand the reaction network are listed in this appendix. A graph transformation rule consists of a subgraph on the left side which can be replaced by the subgraph on the right side when a match of the left-hand side is found in an already existing molecular graph in the network generated so far. Since these graph transformation rules represent a class of reactions, they can be considered as reaction templates.
[…]
Figure 4: Rule 1: Addition of water or ammonia to hydrogen cyanide (nitrile group) where $X\in\{\mathrm{N},\mathrm{O}\}$ to form an imine. The reverse of this rule was used while expanding the hypergraph.
[…]
Figure 5: Rule 2: Addition of water or ammonia to a double bond where $X,Y\in\{\mathrm{N},\mathrm{O}\}$ (in an imine or carbonyl group respectively). The reverse of this rule was also used while expanding the hypergraph.
[…]
Figure 6: Rule 3: 1,2 to 2,3 bond shift of the double bond to form tautomers where $X,Y\in\{\mathrm{N},\mathrm{O}\}$. The reverse of this rule was also used while expanding the hypergraph.
Appendix B List of Molecules in the presented pathways
In this appendix, we supply the molecular structure and the corresponding keys of the vertices (in the form $v_{i}$, where $i$ is some integer) in the pathways that are presented in Section IV. The two tables on the left side depict the molecules in the pathways presented in Figure 2, while the table on the right side depicts the molecules in the pathways presented in Figure 3.
Name Molecular Structure
$v_{0}$chemfigHO-[:30]-[:-30] [:-30]N
$v_{1}$chemfigNH_3
$v_{2}$chemfigH_2O
$v_{3}$chemfigHO-[:30]-[:-30](-[:-90]OH)=[:30]NH
$v_{4}$chemfigHO-[:30]=[:-30](-[:-90]OH)-[:30]NH_2
$v_{5}$chemfigO=[:30]-[:-30](-[:-90]OH)-[:30]NH_2
$v_{6}$chemfigO=[:30]-[:-30]=[:30]NH
$v_{7}$chemfigHO-[:30](-[:90]OH)-[:-30]=[:30]NH
$v_{8}$chemfigHO-[:30](-[:90]OH)=[:-30]-[:30]NH_2
$v_{9}$chemfigHO-[:30](=[:90]O)-[:-30]-[:30]NH_2
Name Molecular Structure
$v_{13}$chemfigHO-[:30]-[:-30](-[:-90]NH_2)=[:30]NH
$v_{14}$chemfigHO-[:30]=[:-30](-[:-90]NH_2)-[:30]NH_2
$v_{15}$chemfigO=[:30]-[:-30](-[:-90]NH_2)-[:30]NH_2
$v_{16}$chemfigHO-[:30](-[:90]OH)-[:-30](-[:-90]NH_2)-[:30]NH_2
$v_{36}$chemfigH_2N-[:30](-[:90]OH)-[:-30](-[:-90]OH)-[:30]NH_2
$v_{52}$chemfigH_2N-[:30](=[:90]O)-[:-30]-[:30]OH
Name Molecular Structure
$v_{0}$chemfigHO-[:30]-[:-30] [:-30]N
$v_{1}$chemfigNH_3
$v_{2}$chemfigH_2O
$v_{3}$chemfigHO-[:30]-[:-30](-[:-90]OH)=[:30]NH
$v_{6}$chemfigHO-[:30]-[:-30](-[:-60]OH)(-[:-120]HO)-[:30]NH_2
$v_{7}$chemfigHO-[:30]-[:-30](=[:-90]O)-[:30]OH
$v_{15}$chemfigHO-[:30]-[:-30](=[:-90]O)-[:30]NH_2
$v_{24}$chemfigHO-[:30]=[:-30](-[:-90]OH)-[:30]NH_2
$v_{33}$chemfigHO-[:30]-[:-30](-[:-90]NH_2)=[:30]NH
$v_{34}$chemfigHO-[:30]-[:-30](-[:-60]NH_2)(-[:-120]HO)-[:30]NH_2
Table 3: The molecular structures corresponding to the the labeled vertices in the pathways in Figure 2 (two left columns) and the pathways in Figure 3 (right column).
Appendix C Energy barriers in the presented pathways
The estimated energy barriers for the reactions in the generated reaction network, which are assigned as weights to the hyperedges during the pathway search and scoring process, are presented in the two tables in this appendix. Table 4 consists of five columns that list the assigned weights for the hyperedges in the pathways shown in Figure 2, while Table 5 contains six columns that list the assigned weights for the hyperedges in the pathways illustrated in Figure 3.
The shaded cells indicate hyperedges that have already been used in a previously enumerated pathway. For instance, all five pathways in Figure 2 share the final two reactions. Consequently, the next four pathways reuse the last two reactions of the first pathway, and the last two rows for the corresponding columns are shaded to reflect this overlap.
Reaction Energy barriers (in kJ mol-1)
count          pathway 1          pathway 2          pathway 3          pathway 4          pathway 5
$e_{1}$ $13.95$ $33.18$ $33.18$ $13.95$ $13.95$
$(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{1},v_{13})$ $(v_{0}+v_{1},v_{13})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{2},v_{3})$
$e_{2}$ $7.67$ $64.17$ $64.17$ $7.67$ $18.67,13.15$
$(v_{3},v_{4})$ $(v_{13},v_{14})$ $(v_{13},v_{14})$ $(v_{3},v_{4})$ $(v_{3},v_{52},v_{4})$
$e_{3}$ $57.69$ $28.40$ $28.40$ $57.69$ $57.69$
$(v_{4},v_{5})$ $(v_{14},v_{15})$ $(v_{14},v_{15})$ $(v_{4},v_{5})$ $(v_{4},v_{5})$
$e_{4}$ $1.67$ $19.86$ $29.35$ $67.36$ $1.67$
$(v_{5},v_{6}+v_{2})$ $(v_{15}+v_{2},v_{16})$ $(v_{15},v_{6}+v_{1})$ $(v_{5}+v_{2},v_{36})$ $(v_{5},v_{6}+v_{2})$
$e_{5}$ $17.14$ $23.79$ $17.14$ $60.84$ $17.14$
$(v_{6}+v_{2},v_{7})$ $(v_{16},v_{7}+v_{1})$ $(v_{6}+v_{2},v_{7})$ $(v_{36},v_{7}+v_{2})$ $(v_{6}+v_{2},v_{7})$
$e_{6}$ $0.19$ $0.19$ $0.19$ $0.19$ $0.19$
$(v_{7},v_{8})$ $(v_{7},v_{8})$ $(v_{7},v_{8})$ $(v_{7},v_{8})$ $(v_{7},v_{8})$
$e_{7}$ $30.42$ $30.42$ $30.42$ $30.42$ $30.42$
$(v_{8},v_{9})$ $(v_{8},v_{9})$ $(v_{8},v_{9})$ $(v_{8},v_{9})$ $(v_{8},v_{9})$
sum $128.74$ $200.02$ $202.85$ $238.13$ $152.90$
Table 4: From the left, the columns list the weights assigned to the hyperedges along with the sum of the weights, for the five pathways in Figure 2.
Reaction Energy barrier
count          pathway 1          pathway 2          pathway 3          pathway 4          pathway 5          pathway 6
$e_{1}$ $13.95$ $13.95$ $13.95$ $33.18$ $13.95$ $33.18$
$(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{1},v_{33})$ $(v_{0}+v_{2},v_{3})$ $(v_{0}+v_{1},v_{3}3)$
$e_{2}$ $7.67$ $48.612$ $80.53$ $48.61$
$(v_{3},v_{24})$ $(v_{33}+v_{2},v_{34})$ $(v_{3}+v_{1},v_{34})$ $(v_{33}+v_{2},v_{34})$
$e_{3}$ $26.61$ $18.68$ $11.76$ $43.64$ $43.64$ $33.08$
$(v_{3}+v_{2},v_{6})$ $(v_{3},v_{15})$ $(v_{24},v_{15})$ $(v_{34},v_{15}+v_{1})$ $(v_{34},v_{15}+v_{1})$ $(v_{34},v_{3}+v_{1})$
$e_{4}$ $0.62$ $0.62$ $0.62$ $0.62$ $0.62$
$(v_{15}+v_{2},v_{6})$ $(v_{15}+v_{2},v_{6})$ $(v_{15}+v_{2},v_{6})$ $(v_{15}+v_{2},v_{6})$ $(v_{3}+v_{2},v_{6})$
$e_{5}$ $74.62$ $74.62$ $74.62$ $74.62$ $74.62$ $74.62$
$(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$ $(v_{6},v_{7}+v_{1})$
sum $118.19$ $107.87$ $108.63$ $200.68$ $213.38$ $219.11$
Table 5: From the left, the columns list the weights assigned to the hyperedges along with the sum of the weights, for the six pathways in Figure 3.
Appendix D Constructed Reaction Network
Figure 7: Hypergraph with $|V|=44$ and $|E|=116$, representing the reaction network considered in this work.